A car's odometer read 22687 km at the start of a trip and 22791 km at end. The trip tool 4.0 hours. What was the car's average speed in km/h and in m/s.


1
Expert's answer
2021-03-09T15:30:32-0500
v=dt,v=\dfrac{d}{t},v=22791 km22687km4 h=26 kmh1000 m1 km1 h3600 s=7.22 ms.v=\dfrac{22791\ km-22687 km}{4\ h}=26\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s}=7.22\ \dfrac{m}{s}.

Answer:

v=26 kmh=7.22 ms.v=26\ \dfrac{km}{h}=7.22\ \dfrac{m}{s}.